OpenStudy (anonymous):

Solve using the quadratic formula. Approximate answers to the nearest tenth. x^2+7x+12=0

OpenStudy (anonymous):

quadratic formula is \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] and standard form ax^2+bx+c=0 so in put the number \[\frac{ -7\pm \sqrt{7^{2}-4(1)(12)} }{ 2(1) }\] \[\frac{ -7\pm \sqrt{49-48} }{ 2 }\] \[\frac{ -7\pm \sqrt{1} }{ 2 }\] \[\frac{ -7\pm 1 }{ 2 }\] \[\frac{ -7+1 }{ 2 }\] and \[\frac{ -7- 1 }{ 2 }\] -6/2 and -8/2 so the answers are x=-3 and -4

OpenStudy (anonymous):

you can check it with factoring since that equation factors into (x+3)(x+4)=0 x+3=0 x=-3 x+4=0 x=-4

OpenStudy (anonymous):

you got a medal

OpenStudy (anonymous):

thank you